OpenStudy (anonymous):

y=In3x, where x=e (equation of line) Don't know how to work with e. Differentiations = 1/x and plugging in value gives 1/e ...How to carry on?

OpenStudy (anonymous):

Wait, what are you trying to find/do?

OpenStudy (anonymous):

Find the equation of the line.

OpenStudy (anonymous):

well the derivative is just the slope at any point. The point is (e), remember (e) is a constant, so 1/e is a real number.

OpenStudy (anonymous):

knowing that, you just need to find the rest of the tangent line: f(x) - f(e) = (1/e)(x - e)

OpenStudy (anonymous):

so just plug (e) back into the original function: ln(3e) = ln3 + ln(e) = ln3 + 1

OpenStudy (anonymous):

the whole thing (i'm assuming) i think should look like: y - (ln3 + 1) = (1/e) (x-e) someone correct me if i'm off.

OpenStudy (anonymous):

The answer to this whole thing is ey=x+eIn3 I just don't know how to get it.

OpenStudy (campbell_st):

1st find y when x = e ln(3e) = ln(3) + 1 since ln(e) = 1 then dy/dx = 3/3x or dy/dx = 1/x to findthe gradient substitute x = e then m = 1/e and point (e, (ln(3)+1)) \[y - (\ln(3)+1) = 1/e(x - e)\] y = x/e - ln(3)

OpenStudy (anonymous):

going off my equation: y - ln3 - 1 = (1/e) (x-e) multiply both sides by (e) ey - eln3 -e = x -e ey = x - e + e + eln3 ey = x + ln3

OpenStudy (anonymous):

hint: y prime=slope. given the x or the x intercept we can find the equation of the line y at 0.

OpenStudy (campbell_st):

for your answer order... they multiplied through by e to get rid of the fraction

OpenStudy (anonymous):

ey = x + eln3* sorry made a typo :)

OpenStudy (anonymous):

Oh. Ok. I need to look at this again. Thanks guys!

OpenStudy (campbell_st):

small problem jerwyn.... you need to substitute x = e to find the value of y ... don't assume